# Hilbert Space Fragmentation and Subspace Scar Time-Crystallinity in Driven Homogeneous Central-Spin Models

### v1

@@ Abstract
We study the stroboscopic non-equilibrium quantum dynamics of periodically kicked Hamiltonians involving homogeneous central-spin interactions. The system exhibits a strong fragmentation of Hilbert space into four-dimensional Floquet-Krylov subspaces, which oscillate between two disjointed two-dimensional subspaces and thus break the discrete time-translation symmetry of the system. Our analytical and numerical analyses reveal that fully polarized states of the satellite spins exhibit fragmentations that are stable against perturbations and have high overlap with Floquet eigenstates of atypically low bipartite entanglement entropy (scar states). We present evidence of robust time-crystalline behavior in the form of a period doubling of the total magnetization of fully polarized satellite spin states that persists over long time scales. We compute non-equilibrium phase diagrams with respect to a magnetic field, coupling terms, and pulse error for various interaction types, including Heisenberg, Ising, XXZ, and XX. We also discuss possible experimental realizations of scar time crystals in color center, quantum dot, and rare-earth ion platforms.
@@ I Introduction
In this paper, we show that Hilbert space fragmentation and scar time-crystallinity can occur in a driven, homogeneous central-spin model. We consider a system with homogeneous interactions (of Heisenberg, Ising, XXZ, or XX type) subject to periodic $\pi$-pulse driving. Within a given symmetry sector of the satellite spin Hilbert space, we find that the stroboscopic dynamics are constrained to a four-dimensional Floquet-Krylov subspace, leading to a fragmentation of Hilbert space. The Floquet-Krylov subspace oscillates between two disjointed regions, breaking the discrete time-translation symmetry of the system. The fully polarized satellite states exhibit a more robust fragmentation compared to other initial states, and they have a high overlap with Floquet scar states—the Floquet eigenstates that have low bipartite entanglement entropy. Furthermore, we diagnose the scar time-crystallinity using total satellite spin magnetization. We note that, in this paper, scar time-crystalline behavior is identified based on the long-lived (but possibly finite) period doubling that requires interactions and is robust to pulse errors for a finite number of satellite spins—we do not consider the thermodynamic limit, where fundamental conceptual questions about the nature of non-equilibrium phases remain unclear [34]. We show that the stability of the fragmentation and time-crystalline behavior against pulse errors depends on the initial state, the nature of the interaction, and the magnetic field on the central spin. We find that the stability is the most prevalent for Ising interactions, followed by XXZ, Heisenberg, and XX interactions, respectively. Numerical simulations reveal a robust subharmonicity (period doubling) in the satellite magnetization for fully polarized satellite states. We obtain non-equilibrium phase diagrams with respect to pulse error, interaction strength, and magnetic field for different types of interactions.
We have organized the paper as follows. In Sec. II, we introduce the homogeneous XXZ central-spin model driven by periodic $\pi$ pulses and study the relevant symmetry of the Floquet Hamiltonian. We also study the Hilbert space fragmentation and its stability. We further calculate the bipartite entanglement entropy of the satellite part of a Floquet eigenstate and its overlap with fully polarized satellite spin states. In Sec. III, we focus on scar time-crystallinity in the system. To diagnose the time-crystalline behavior, we calculate the satellite magnetization numerically. Period doubling of the satellite magnetization is shown by plotting it with respect to various parameters like the magnetic field, coupling terms, and pulse error. In Sec. IV, we also discuss the experimental realization of time crystals in color center, rare-earth ion, and quantum dot platforms. We conclude in Sec. V. Some technical details are given in the appendices.
@@ II.1 Model Hamiltonian
We consider a central-spin model consisting of one central spin coupled to $N$ satellite spins via XXZ-type interactions with uniform coupling strengths (see Fig. 1(a)). In the context of color centers, rare earth ions, or quantum dots, the central spin is an electronic spin that couples to nuclear (satellite) spins via hyperfine interactions. The form of such interactions can vary from Heisenberg to Ising depending on the details of the system in question. The XXZ model encompasses all of these possibilities. We can write the Hamiltonian as
$H_{0}$ $=\sum_{p=1}^{N}\Big{[}A_{xy}(\mathrm{I}_{p}^{x}S_{0}^{x}+\mathrm{I}_{p}^{y}S_{0}^{y})+A_{z}\mathrm{I}_{p}^{z}S_{0}^{z}+B^{n}_{\mu}\mathrm{I}_{p}^{\mu}\Big{]}$
@@
The $z$-component of the total spin of the system is a conserved quantity, $\big{[}H,S_{0}^{z}+\sum_{p=1}^{N}\mathrm{I}_{p}^{z}\big{]}=0$, when the Zeeman fields are in the $z$ direction. The magnitude of the total satellite spin is also conserved, $\big{[}H,\mathrm{I}_{t}^{2}\big{]}=0$, where
@@
In the context of stroboscopic dynamics, where we focus only on the evolution at a discrete set of times corresponding to the moments immediately after each pulse, if an operator commutes with the Floquet Hamiltonian, then it corresponds to a symmetry of the system. We note that $\big{[}H_{p},S_{0}^{z}+\mathrm{I}_{t}^{z}\big{]}\neq 0\Rightarrow\big{[}H_{p},H_{F}\big{]}\neq 0$, and so therefore $S_{0}^{z}+\mathrm{I}_{t}^{z}$ is no longer conserved when the system is driven. On the other hand, $[\mathrm{I}_{t}^{2},\mathrm{I}_{t}^{\mu}]=0$ implies that $[\mathrm{I}_{t}^{2},H_{0}]=0$ and $[\mathrm{I}_{t}^{2},H_{\text{p}}]=0$, which implies that
$[H_{F},\mathrm{I}_{t}^{2}]=0\text{ and }[U_{F},\mathrm{I}_{t}^{2}]=0,$ (6)
@@ II.2 Floquet-Krylov subspaces and Hilbert space fragmentation
This fragmentation of $U_{F}$ is depicted in Fig. 1(c) (blue boxes). The evolution in each 4D Floquet-Krylov subspace consists of oscillations between two disjointed 2D subspaces. For example, the initial state $\mathinner{|{m_{j}\uparrow}\rangle}$ first evolves in the 2D subspace spanned by $\mathinner{|{m_{j}\uparrow}\rangle}$ and $\mathinner{|{(m_{j}+1)\downarrow}\rangle}$. The first $\pi$ pulse then maps this into the separate 2D subspace spanned by $\mathinner{|{-m_{j}\downarrow}\rangle}$ and $\mathinner{|{-(m_{j}+1)\uparrow}\rangle}$, after which the spin-spin interactions rotate the state in this subspace. Each subsequent $\pi$ pulse flips the state from one 2D subspace to the other, such that each subspace is visited every two driving periods. This $2T$ oscillatory behavior breaks the discrete time-translation symmetry of the time-periodic Hamiltonian.
We see that even if we increase the number of nuclei $N$, the Floquet-Krylov subspace remains 4D, while the dimension of the largest symmetry sector corresponding to $\mathrm{I}_{t}^{2}$ increases linearly as $N+1$, and the full Hilbert space dimension increases exponentially as $2^{N}$. Thus, the ratio of the dimension of the largest stroboscopic dynamical subspace to the dimension of the full Hilbert space (or symmetry sector) goes to zero in the large $N$ limit. This type of fragmentation is the Floquet version of strong HSF [28] and leads to a non-ergodic behavior of the system. Note that this fragmentation holds for all values of the spin-spin couplings and Zeeman field strengths in the absence of pulse errors. Next, we show that the $2T$ oscillatory behavior is stable against small perturbations $\theta$ to the pulse angle, indicating the existence of time-crystalline behavior in the system without any random coupling disorder.
@@ II.3 Evidence of scar states from entanglement entropy
Results for the central-spin model with Heisenberg interactions are shown in Fig. 3. In particular, the figure shows the entanglement entropy of each Floquet eigenstate and its overlap with both fully polarized satellite states with ($\theta=0.2\pi$) and without ($\theta=0$) pulse errors. We observe that the orange-colored dots (eigenstates) possess atypically low bipartite entanglement entropy and have high overlap with the satellite states |±$\mathinner{|{\pm J}\rangle}$. Therefore, for both values of $\theta$, the satellite states |±$\mathinner{|{\pm J}\rangle}$ are closely associated to scar states (Floquet eigenstates with atypically low bipartite entanglement entropy) of the system. Additionally, in Appendix C, we show the entanglement entropy of Floquet eigenstates of the driven XXZ central-spin system. There too, fully polarized satellite states exhibit the characteristics of scar states.
@@
III Scar time-crystals
Next, we show that periodically driven central-spin models can exhibit scar time-crystalline behavior in certain parameter regimes. We define the scar time-crystal as a non-equilibrium phase that shows a robust subharmonic response only for a few initial states associated with the weak breaking of ergodicity. To diagnose the time-crystalline behavior, we need to calculate the average satellite spin magnetization at integer multiples of the driving period. For an initial state $\mathinner{|{\Psi(0)}\rangle}$, we have
where $\mathinner{|{\Psi(nT)}\rangle}=U(nT)\mathinner{|{\Psi(0)}\rangle}$, $N$ is the number of satellite spins, and $\mathrm{I}_{t}^{z}=\sum_{p=1}^{N}I_{p}^{z}$ as before. When $\mathinner{\langle{\mathrm{I}_{t}^{z}((n+1)T)}\rangle}\neq\mathinner{\langle{\mathrm{I}_{t}^{z}(nT)}\rangle}$ for large $n$, then the average magnetization explicitly breaks the discrete time-translation symmetry of the system. This is a signature of the time-crystalline behavior of the system.
In this section, we first show that only a few initial states that exhibit stable HSF and have high overlap with Floquet scar states exhibit a robust period doubling of the total satellite spin magnetization. Therefore the system shows scar time-crystalline behavior. We then consider the four initial states ({|±$\{\mathinner{|{\pm J\sigma}\rangle}\}$) and investigate the time crystallinity with respect to coupling strength, Zeeman energy of the central spin, and pulse error.
III.1 Scar time-crystallinity in Ising, XX, XXX, and XXZ central-spin models

### v2

@@ Abstract
We study the stroboscopic non-equilibrium quantum dynamics of periodically kicked Hamiltonians involving homogeneous central-spin interactions. The system exhibits a strong fragmentation of Hilbert space into four-dimensional Floquet-Krylov subspaces, which oscillate between two disjointed two-dimensional subspaces and thus break the discrete time-translation symmetry of the system. Our analytical and numerical analyses reveal that fully polarized states of the satellite spins exhibit fragmentations that are stable against perturbations and have high overlap with Floquet eigenstates of atypically low bipartite entanglement entropy (scar states). Motivated by the breaking of discrete time translation symmetry by Floquet-Krylov subspaces, we introduce a novel type of time crystal that we call a “subspace time crystal”. We present evidence of robust time-crystalline behavior in the form of a period doubling of the total magnetization of fully polarized satellite spin states that persists over long time scales. We compute non-equilibrium phase diagrams with respect to a magnetic field, coupling terms, and pulse error for various interaction types, including Heisenberg, Ising, XXZ, and XX. We also discuss possible experimental realizations of scar time crystals in color center, quantum dot, and rare-earth ion platforms.
@@ I Introduction
In this paper, we show that Hilbert space fragmentation and scar time-crystallinity can occur in a driven, homogeneous central-spin model. We consider a system with homogeneous interactions (of Heisenberg, Ising, XXZ, or XX type) subject to periodic $\pi$-pulse driving. Within a given symmetry sector of the satellite spin Hilbert space, we find that the stroboscopic dynamics are constrained to a four-dimensional Floquet-Krylov subspace, leading to a fragmentation of Hilbert space. The Floquet-Krylov subspace oscillates between two disjointed regions, breaking the discrete time-translation symmetry of the system. The fully polarized satellite states exhibit a more robust fragmentation compared to other initial states, and they have a high overlap with Floquet scar states—the Floquet eigenstates that have low bipartite entanglement entropy. Furthermore, we define “subspace time crystals” using the concept of Floquet Krylov subspaces. We diagnose the scar time-crystallinity using total satellite spin magnetization. We note that, in this paper, scar time-crystalline behavior is identified based on the long-lived (but possibly finite) period doubling that requires interactions and is robust to pulse errors for a finite number of satellite spins—we do not consider the thermodynamic limit, where fundamental conceptual questions about the nature of non-equilibrium phases remain unclear [34]. We show that the stability of the fragmentation and time-crystalline behavior against pulse errors depends on the initial state, the nature of the interaction, and the magnetic field on the central spin. We find that the stability is the most prevalent for Ising interactions, followed by XXZ, Heisenberg, and XX interactions, respectively. Numerical simulations reveal a robust subharmonicity (period doubling) in the satellite magnetization for fully polarized satellite states. We obtain non-equilibrium phase diagrams with respect to pulse error, interaction strength, and magnetic field for different types of interactions.
We have organized the paper as follows. In Sec. II, we introduce the homogeneous XXZ central-spin model driven by periodic $\pi$ pulses and study the relevant symmetry of the Floquet Hamiltonian. We also study the Hilbert space fragmentation and its stability. We further calculate the bipartite entanglement entropy of the satellite part of a Floquet eigenstate and its overlap with fully polarized satellite spin states. In Sec. III, we formally define subspace time crystals and focus on scar time-crystallinity in the system. To diagnose the time-crystalline behavior, we calculate the satellite magnetization numerically. Period doubling of the satellite magnetization is shown by plotting it with respect to various parameters like the magnetic field, coupling terms, and pulse error. In Sec. IV, we also discuss the experimental realization of time crystals in color center, rare-earth ion, and quantum dot platforms. We conclude in Sec. V. Some technical details are given in the appendices.
@@ II.1 Model Hamiltonian
We consider a central-spin model consisting of one central spin coupled to $N$ satellite spins via XXZ-type interactions with uniform coupling strengths (see Fig. 1(a)). In the context of color centers, rare earth ions, or quantum dots, the central spin is an electronic spin that couples to nuclear (satellite) spins via hyperfine interactions. The form of such interactions can vary from Heisenberg to Ising depending on the details of the system in question. A homogeneous central-spin system can effectively represent a dense ensemble of nuclei in a quantum dot. Such systems have been studied in experiments, as discussed in [45]. Additionally, homogeneous Heisenberg couplings can describe a GaAs quantum dot realistically at short times and low magnetic fields [46, 47, 48]. The XXZ model encompasses all of these possibilities. We can write the Hamiltonian as
$H_{0}$ $=\sum_{p=1}^{N}\Big{[}A_{xy}(\mathrm{I}_{p}^{x}S_{0}^{x}+\mathrm{I}_{p}^{y}S_{0}^{y})+A_{z}\mathrm{I}_{p}^{z}S_{0}^{z}+B^{S}_{\mu}\mathrm{I}_{p}^{\mu}\Big{]}$
@@
The $z$-component of the total spin of the system is a conserved quantity, $\big{[}H_{0},S_{0}^{z}+\sum_{p=1}^{N}\mathrm{I}_{p}^{z}\big{]}=0$, when the Zeeman fields are in the $z$ direction. The magnitude of the total satellite spin is also conserved, $\big{[}H_{0},\mathrm{I}_{t}^{2}\big{]}=0$, where
@@
In the context of stroboscopic dynamics, where we focus only on the evolution at a discrete set of times corresponding to the moments immediately after each pulse, if an operator commutes with the Floquet Hamiltonian, then it corresponds to a symmetry of the system. We note that $\big{[}H_{p},S_{0}^{z}+\mathrm{I}_{t}^{z}\big{]}\neq 0\Rightarrow\big{[}S_{0}^{z}+\mathrm{I}_{t}^{z},H_{F}(\theta)\big{]}\neq 0$, and therefore $S_{0}^{z}+\mathrm{I}_{t}^{z}$ is no longer conserved when the system is driven. On the other hand, $[\mathrm{I}_{t}^{2},\mathrm{I}_{t}^{\mu}]=0$ implies that $[\mathrm{I}_{t}^{2},H_{0}]=0$ and $[\mathrm{I}_{t}^{2},H_{\text{p}}]=0$, which implies that
$[H_{F}(\theta),\mathrm{I}_{t}^{2}]=0\text{ and }[U_{F}(\theta),\mathrm{I}_{t}^{2}]=0,$ (6)
@@ II.2 Floquet-Krylov subspaces and Hilbert space fragmentation
This fragmentation of $U_{F}$ is depicted in Fig. 1(c) (blue boxes). The evolution in each 4D Floquet-Krylov subspace consists of oscillations between two disjointed 2D subspaces. For example, the initial state $\mathinner{|{m_{j}\uparrow}\rangle}$ first evolves in the 2D subspace spanned by $\mathinner{|{m_{j}\uparrow}\rangle}$ and $\mathinner{|{(m_{j}+1)\downarrow}\rangle}$. The first $\pi$ pulse then maps this into the separate 2D subspace spanned by $\mathinner{|{-m_{j}\downarrow}\rangle}$ and $\mathinner{|{-(m_{j}+1)\uparrow}\rangle}$, after which the spin-spin interactions rotate the state in this subspace. Each subsequent $\pi$ pulse flips the state from one 2D subspace to the other, such that each subspace is visited every two driving periods. This $2T$ oscillatory behavior breaks the discrete time-translation symmetry of the time-periodic Hamiltonian. To classify this phenomenon more formally, we introduce the notion of subspace time-crystallinity for general time-periodic Hamiltonians in Sec. III.1.
We see that even if we increase the number of satellite spins $N$, the Floquet-Krylov subspace remains 4D, while the dimension of the largest symmetry sector corresponding to $\mathrm{I}_{t}^{2}$ increases linearly as $N+1$, and the full Hilbert space dimension increases exponentially as $2^{N}$. Thus, the ratio of the dimension of the largest stroboscopic dynamical subspace to the dimension of the full Hilbert space (or symmetry sector) goes to zero in the large $N$ limit. This type of fragmentation is the Floquet version of strong HSF [28] and leads to a non-ergodic behavior of the system. Note that this fragmentation holds for all values of the spin-spin couplings and Zeeman field strengths in the absence of pulse errors. Next, we show that the $2T$ oscillatory behavior is stable against small perturbations $\theta$ to the pulse angle, indicating the existence of time-crystalline behavior in the system without any random coupling disorder.
@@ II.3 Evidence of scar states from entanglement entropy
Results for the central-spin model with Heisenberg interactions are shown in Fig. 3. In particular, the figure shows the entanglement entropy of each Floquet eigenstate and its overlap with both fully polarized satellite states with ($\theta=0.2\pi$) and without ($\theta=0$) pulse errors. We observe that the orange-colored (circled) dots (eigenstates) possess atypically low bipartite entanglement entropy and have high overlap with the satellite states |±$\mathinner{|{\pm J}\rangle}$. Therefore, for both values of $\theta$, the satellite states |±$\mathinner{|{\pm J}\rangle}$ are closely associated to scar states (Floquet eigenstates with atypically low bipartite entanglement entropy) of the system. Additionally, in Appendix C, we show the entanglement entropy of Floquet eigenstates of the driven XXZ central-spin system. There too, fully polarized satellite states exhibit the characteristics of scar states.
III Scar time-crystals
Next, we show that periodically driven central-spin models can exhibit (subspace) scar time-crystalline behavior in certain parameter regimes. In this section, we first define subspace time crystals using the concept of Floquet Krylov subspaces for a general time-periodic Hamiltonian and discuss signatures of time-crystallinity in our model. We then show that only a few initial states that exhibit stable HSF and have high overlap with Floquet scar states exhibit a robust period doubling of the total satellite spin magnetization. Therefore the system shows scar time-crystalline behavior. We then consider the four initial states ({|±$\{\mathinner{|{\pm J\sigma}\rangle}\}$) and investigate the time crystallinity with respect to coupling strength, Zeeman energy of the central spin, and pulse error.
III.1 Subspace time crystal
We consider a time-periodic Hamiltonian $H(t+T)=H(t)$ in a Hilbert space $\mathcal{H}$. We decompose this Hilbert space in terms of Floquet Krylov subspaces as $\mathcal{H}=\oplus_{i}\mathcal{K}_{Fi}$, where $\mathcal{K}_{Fi}=\mathcal{K}_{F}(U_{F},\mathinner{|{\psi_{i}}\rangle})$ is defined in Eq. (II.2). Each $\mathcal{K}_{Fi}$ consists of disjointed subspaces such that $\mathcal{K}_{F}=\oplus_{q=1}^{r}\mathcal{K}_{F}^{q}$, where $r\in\mathbb{N}$. These disjointed subspaces are obtained from the spans of states reached after integer multiples of the drive period. For an experimentally accessible quantum state (used in defining the Floquet Krylov subspace) and $n\in\mathbb{N}$, if we find that
$U_{F}^{n}\mathinner{|{\psi}\rangle}\in\mathcal{K}_{F}^{q},\text{but }U_{F}^{n+r_{0}}\mathinner{|{\psi}\rangle}\notin\mathcal{K}_{F}^{q}\ \forall\ r_{0}<r\in\mathbb{N},\text{ and}$
$\hskip 71.13188ptU_{F}^{n+r}\mathinner{|{\psi}\rangle}=\mathinner{|{\psi^{\prime}}\rangle}\in\mathcal{K}_{F}^{q},$ (15)
then the system periodically returns to the same subspace with period $rT$ and thus breaks the discrete time translation symmetry. When this breaking remains stable under small local perturbations, we say the system exhibits subspace time-crystallinity. If all subspaces show the robustness against the perturbations then we have “full (MBL type) subspace time-crystallinity”. But if the robustness exists only for a few states, then we have “subspace scar time-crystallinity”. We note that for the existence of a subspace TC, some Floquet-Krylov subspaces must contain at least two disjointed subspaces ($r>1$). Additionally, we can recast a conventional (MBL) TC in terms of two dimensional Floquet-Krylov subspaces containing two 1D disjointed subspaces.
To diagnose a subspace TC, we can calculate the probability (fidelity) to return to the subspace as $\mathbb{F}^{n}_{\Sigma}=\sum_{i}|\mathinner{\langle{\phi_{i}}|{U_{F}^{n}}|{\psi}\rangle}|^{2}$, where $\{\mathinner{|{\phi_{i}}\rangle}\}$ are orthonormal bases of the disjointed subspace containing $\mathinner{|{\psi}\rangle}$. When the system periodically returns to the initial subspace after $r$ drive periods, then we obtain $\mathbb{F}^{n}_{\Sigma}=\delta_{n(jr)}$, $j\in\mathbb{N}$. Other diagnostic probes such as an expectation value of an observable in time may not be useful as the evolved state may be a superposition of eigenstates of the observable. However, in some cases (in our model too), some observable (net magnetization) works well as the dimension of the disjointed subspace is small (two-dimensional in our model).
We can further outline a general framework to identify subspace time-crystallinity whenever there is Hilbert space fragmentation for a time-independent Hamiltonian. Consider a time-independent Hamiltonian $H_{0}$ that exhibits HSF, so that $\mathcal{H}=\oplus_{i}\mathcal{K}_{i}$, where $\mathcal{K}_{i}=\mathcal{K}(H_{0},\mathinner{|{\psi_{i}}\rangle})$ defined in Eq. (II.2). We choose two non-ergodic subspaces (say $\mathcal{K}_{1},\mathcal{K}_{2}$) such that we can connect them using some local unitary operator ($U_{L}$), i.e., if a state $\mathinner{|{\psi}\rangle}\in\mathcal{K}_{1}$ then $U_{L}\mathinner{|{\psi}\rangle}\in\mathcal{K}_{2}$ and vice-versa. Thus, for Floquet operator $U_{F}=U_{L}e^{-iH_{0}T}$, we obtain a subspace TC ($\mathcal{K}_{F}(U_{F},\mathinner{|{\psi}\rangle})=\mathcal{K}_{1}\oplus\mathcal{K}_{2}$) if it also shows robustness against small local perturbations.
Now we focus on our central spin model. We define the (subspace) scar time-crystal as a non-equilibrium phase that shows a robust subharmonic response only for a few initial states associated with the weak breaking of ergodicity. To diagnose the time-crystalline behavior, we need to calculate the average satellite spin magnetization at integer multiples of the driving period. For an initial state $\mathinner{|{\Psi(0)}\rangle}$, we have
where $\mathinner{|{\Psi(nT)}\rangle}=U(nT)\mathinner{|{\Psi(0)}\rangle}$, $N$ is the number of satellite spins, and $\mathrm{I}_{t}^{z}=\sum_{p=1}^{N}\mathrm{I}_{p}^{z}$ as before. When $\mathinner{\langle{\mathrm{I}_{t}^{z}((n+1)T)}\rangle}\neq\mathinner{\langle{\mathrm{I}_{t}^{z}(nT)}\rangle}$ for large $n$, then the average magnetization explicitly breaks the discrete time-translation symmetry of the system. This is a signature of the time-crystalline behavior of the system. We note that for a subspace TC, ideally we should examine the return probability ($\mathbb{F}^{n}_{\Sigma}$) as discussed above. However, because the disjointed subspaces (Krylov subspaces of $H_{0}$) are maximally two-dimensional and eigenstates of $\mathrm{I}_{t}^{z}$, the net satellite magnetization works well. Henceforth in this text, we use the term “scar TC” to refer to what is in fact a subspace scar TC for convenience. We will discuss the nature of time-crystallinity corresponding to different interactions and initial states in Sec. III.4.
@@ III.1 Subspace time crystal
III.2 Scar time-crystallinity in Ising, XX, XXX, and XXZ central-spin models